Here are book- and market-value balance sheets of the United Frypan Company:

Book-Value Balance Sheet Net working capital $ 50 Long-term assets 50 Total:100 Equity $30 Debt $70 Total:100

Market-Value Balance Sheet Net working capital $ 50 Long-term assets 200 Total:250 Equity $180 Debt $70 Total: 250

Assume that MM’s theory holds except for taxes. There is no growth, and the $70 of debt is expected to be permanent. Assume a 32% corporate tax rate.

a. How much of the firm's market value is accounted for by the debt-generated tax shield?

b. What is United Frypan’s after-tax WACC if rDebt = 7.4% and rEquity = 15.6%?

c. Now suppose that Congress passes a law that eliminates the deductibility of interest for tax purposes after a grace period of 5 years. What will be the new value of the firm, other things equal? Assume a borrowing rate of 7.4%.
Answer #1

a. Market Value accounted by debt generated tax shield = Debt* Tax rate = 70*32% = 22.40

b. After Tax WACC = Weight of Debt * Cost of Debt*(1-Tax rate)+ Weight of equity* Cost of Equity = 70/(70+180)*7.4%*(1-32%)+180/250*15.6% = 12.64%

c. Debt Tax Shield after 5 years = Debt* Tax = 70*32% = 22.40
PV of Tax shield = 22.50/(1+7.4%)^5 = 15.75
Value of firm = Value of Equity + Value of Debt + PV of Debt Tax Shield = 180 + 70 + 15.75 = 265.75

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well